UT September 08, 2019 - Restart Night #7


TELESCOPE CLOSED DUE TO WATER ISSUES/STORM


Plan Outline:

  • LUCI
    • AGW-1 Transform
      • No need to update PCS, rotator ZP looks good after astrometric mask test
    • IF spectroscopic modes are released for LUCIs
      • G200 science-like OB
      • G210 science-like OB
  • Test patch to GCS/PCS to address dithering NS guiding
    • LUCI
    • MODS
  • DX M3 alignment
  • LBC
    • Standard photometric fields - requires clear skies (any seeing) - required for ZP calibrations
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
    • Twilight Flats all filters
    • LBC sub-framing (test OBs needed)
  • MODS
    • Test NS guiding with dithering


Summary


23:25 Confirmed MCE/ROE for LUCI1 was powered on after todays work. Initialized LUCI's which looked good and took test darks. A little bit noisy which improved. Expected after the ROE powered off for extended time. Lookes fine by 23:54. Brought up the RTD on the obs machin through /lbt/lucifer/lcsp/bin/startRTD.sh #. The images were not autopopulating. The readout# directory empty so images manually opened through newdata.

-- BarryRothberg - 07 Sep 2019
Topic revision: r3 - 08 Sep 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback